# Detecting State Changes in Functional Neuronal Connectivity using Factorial Switching Linear Dynamical Systems

### v1

@@ 1 Introduction
The brain receives information from the surrounding world before reacting correspondingly and systematically. As a complex system, it acts differently across time and locations. Thus, analytical tools that can detect correlations in spatial and temporal activity are required to study the brain’s activity. The dynamic networks that underlie these brain processes at the microscale can be studied using microelectrode array (MEA) recordings of neural signals across a whole population of neural cells [27]. While MEA recordings do not allow us to target specific neurons, neural signals, highly correlated in time, from neurons near multiple electrodes can reveal the spatial-temporal features of brain network function. Understanding this functional connectivity in vitro is important for understanding microscale brain function and development in vivo [10].
@@
Previous studies of microscale functional neuronal networks have focused on the network topology [5, 26, 28]. These studies estimate a functional graph by considering pair-wise correlations between electrodes over a period of time. Graph theoretical metrics are then used to compare the topology of the functional networks including the spatial arrangement of nodes [27]. These approaches have two important limitations. Firstly, they estimate a single pattern of functional connectivity based on minutes of MEA recording. This ignores the fact that brain networks are adaptable on much faster time-scales and precludes study of such short-term changes in connectivity. Secondly, they rely on pair-wise comparison of nodal activity to determine functional connectivity, which may not capture patterns of activity observed in subnetworks detected by more than two electrodes.
@@
Other works have used discrete state-space models to capture evolving functional connectivity. These approaches capture the idea that a dynamical system—in our case, temporal MEA recordings—can be explained in terms of a small number of underlying behavioral patterns. In the simplest state-space model, hidden Markov models (HMMs) model neuronal activity as arising from one of a small number of underlying activity patterns [24]. Switching linear dynamical systems (SLDS) extend this by characterizing each activity pattern using a linear dynamical system [25]. Extensions add additional structure between the latent states, allowing for more flexible dynamics [16, 20, 6]. These approaches, however, make the unrealistic assumption that neuronal activity can be partitioned into a small number of a priori independent activity patterns. They lack the capacity to detect multiple network patterns simultaneously, as occurs in complex brain networks.
@@
New methods are needed to identify changes in functional connectivity across seconds to minutes during MEA recordings, without making unrealistic assumptions about the nature of connectivity changes. In this work, we posit that neuronal activity can be decomposed into multiple subnetworks that can be activated jointly or separately. This allows us to capture overlapping patterns of behavior without the need to create a single state for each combination of active subnetworks. Factorial Switching Linear Dynamical Systems (FSLDS) replaces the single discrete latent state model at the heart of SLDS-based models with a factorial latent state model [FHMM, 8], modeling MEA recordings in terms of latent features (or subnetworks) that can be turned on or off independently. Active subnetworks evolve according to a linear dynamical system. To facilitate inference, we formulate our latent state model using a continuous relaxation of the factorial hidden Markov model [18] and perform inference using autoencoded variational Bayes [AEVB, 12]. We first explore properties of FSLDS on simulated data, validating its ability to recover ground-truth additive latent structure. Second, we show how it can uncover interpretable latent structure in MEA recordings. Third, we show how it reveals a developmental increase in the number of latent features (or subnetworks) in human induced pluripotent stem cell (iPSC)-derived neuronal networks in vitro.
@@ 2 Background
Our model is a factorial state-space model that extends existing time-series models by allowing multiple latent factors to be active at once, using a continuous relaxation of a factorial hidden Markov model [FHMM, 8] to control which combination of latent factors are active. This leads to a more interpretable model for temporally evolving neuronal functional connectivity, where we assume that observations are generated from an additive combination of multiple subnetworks, that may not all activate synchronously. In this section, we review existing work in state space models for time series data, with a particular focus on models appropriate for modeling MEA recordings, and discuss inference approaches appropriate for complex discrete state-space models. We then review the FHMM in Section 2.3, which will form the backbone of our approach.
@@ 2.1 Modeling functional connectivity using discrete state space models
Discrete state-space models such as hidden Markov models  [HMMs, 7] provide an interpretable model for time-series data, where we assume that each time-point belongs to one of a finite number of states. Observations are assumed iid given a state-specific emission-distribution. HMMs have been used to uncover latent structure in neuronal activity [3, 32, 15]. Switching dynamical systems replace the emission model with a dynamical system. The simplest example of this is a switching linear dynamical system  [SLDS, 1], where a Markovian discrete state process switches between autoregressive models with state-specific parameters, which can lead to a more flexible model for MEA recordings [25]. Building on SLDS, more sophisticated models have been developed to handle more complex scenarios, such as recurrent switching linear dynamical system  [rSLDS, 16] and tree-structured rSLDS [20]. These models introduce recurrent dependency structures, allowing memory between consecutive time points. Additionally, when linear assumptions prove insufficient, nonlinear dynamical system methods have been proposed [4, 29][21] introduces a variational RNN model with attention to incorporate both nonlinear dynamics and non-Markovian switches.
2.2 Bayesian inference for discrete state space models
The posterior distributions over the parameters of discrete state space models are often learned through Markov chain Monte Carlo (MCMC) using forward filtering backward sampling [31, 9, 22]. However, as model complexity increases, MCMC methods can suffer from slow mixing and high computational cost. Recent advances in variational inference provide a more computationally efficient alternative.
Auto-Encoding Variational Bayes (AEVB) [12] offers a flexible “black box” variational inference framework by using neural networks to specify a variational approximation and using gradient descent to optimize a sample-based estimate of the ELBO. However, gradient-based optimization methods cannot directly learn discrete variables, such as those present in switching dynamical systems. The REINFORCE algorithm [30] uses a Monte Carlo estimate of the gradient; however this estimate suffers from high variance [23]. Other approaches combine variational inference with sequential Monte Carlo [17, 19, 14, 13].
An alternative approach is to sidestep the use of discrete distributions altogether. The Concrete, or Gumbel-Softmax, distribution [11, 18],
$p_{\alpha,\phi}(x)=(K-1)!\phi^{K-1}\prod_{k=1}^{K}\left(\frac{\alpha_{k}x_{k}^{-\phi-1}}{\sum_{j=1}^{K}\alpha_{j}x_{j}^{-\phi}}\right)$ (1)
is a continuous distribution over the probability simplex that can be used as an approximation to a Categorical($\alpha_{1},\dots,\alpha_{K}$) distribution. The temperature parameter $\phi$ controls the smoothness of the distribution; by annealing the temperature $\phi$, the posterior distribution of the continuous approximation asymptotically converges to the discrete Categorical distribution as $\phi\rightarrow 0$. Since the Concrete($\alpha,\phi$) distribution is continuous, we can use it directly in variational inference as an alternative to the Categorical distribution.
2.3 Factorial Hidden Markov Models
[…]
@@ 2.3 Factorial Hidden Markov Models
HMMs (and the more complex switching dynamical systems described in Section 2.1) assume that, at any given time point, observations are generated according to a single latent emission model. This may be unrealistic in complex time series such as MEA recordings.
In practice, two timepoints may share some characteristics, but not others. This observation motivates the factorial hidden Markov model [FHMM, 8], where multiple hidden factors $\{h_{t}^{(k)}\}_{k=1}^{K}$ concurrently influence the observed dynamics at each time point $y_{t}$.
@@
3.1 ConcreteFHMM
The factorial nature of the FHMM leads to an exponentially growing parameter space. Exploration of this discrete state space is challenging, and traditional MCMC methods are prone to difficult and slow mixing [8]. As we discussed in Section 2.2, variational inference provides an appealing alternative to MCMC, allowing us to exploit optimization techniques to learn an approximate posterior. However, as we saw, a naive implementation of variational inference methods does not allow for discrete latent states, since gradient descent is inherently designed for continuous variables.
To overcome this limitation, we introduce a continuous variant of the FHMM by relaxing the discrete latent state indicators into continuous variables using the Concrete distribution (Equation 1). We assume $h_{t}^{(k)}\in[0,1]$ and an additive emission model $f_{\theta}$, so that $h_{t}^{(k)}$ can be seen as turning on or off the $k$th latent feature.
$h_{t}^{(k)}|h_{t-1}^{(k)}=$ $\text{Concrete}\left(f_{p}\left(h_{t-1}^{(k)},y_{t-1}\right),\phi\right)$ (3)
@@ 3.1 ConcreteFHMM
In a FHMM, each latent factor is associated with a transition matrix that describes the probability of transitioning between states. In the ConcreteFHMM, we no longer have discrete state indicators (as $h_{t}^{(k)}$ is continuous). Since we cannot use these continuous indicators to directly index a row of the transition matrix, we introduce a function $f_{p}$ to specify the distribution over the value of the $k$th latent factor at time $t$. In our experiments, we parametrize $f_{p}$ using a recurrent neural network.
@@ 3.2 Factorial switching linear dynamical systems
As we have seen, HMMs, SLDSs, and their variants assume each time point corresponds to a single latent state. In practice, this might be unrealistic. For example, in developing brains, we expect new subnetworks to appear in addition to previous structure, rather than fully replacing the previous connectivity pattern. Further, we do not expect a subnetwork to have constant activity—instead, we want to disentangle changes in activity amplitude from changes in neuron connectivity. To achieve this, we separate these processes by first determining whether a feature is active using a ConcreteFHMM, and then modeling the value of the state when it is active using a linear dynamical system. The generative model is
@@
$h_{t}^{(k)}\mid h_{t-1}^{(k)}$ $\sim\text{Concrete}\left(f_{p}\left(h_{t-1}^{(k)},\,y_{t-1}\right)\right),$ $k=1,\dots,K$
$z_{t}^{(k)}\mid z_{t-1}^{(k)}$ $\sim\text{Normal}\left(Az_{t-1}^{(k)},\,\sigma_{p}^{2}\mathbb{I}\right),$ $k=0,\dots,K.$
The level of activity of the $k$th subnetwork is governed by two components: an (approximately) discrete state $h_{t}^{(k)}$, and a continuous state $z_{t}^{(k)}$. The approximately discrete state $h_{t}^{(k)}$ controls whether a subnetwork is currently contributing to the rate, and aids in interpretability of the latent structure. Meanwhile the continuous states $z_{t}^{(k)}$ control variation in the amplitude of the activity. In the context of MEA recordings, we can think of $h_{t}^{(k)}$ as indicating whether a subnetwork is currently active, and $\exp\left\{z_{t}^{(k)}\right\}$ as controlling variation in its activity while it is active. We summarize FSLDS in Figure 2.
The latent feature indicators $h_{t}^{(k)}$ follow a ConcreteHMM, as described in Section 3.1. The choice of a ConcreteFHMM is primarily due to computational practicality; as we will see in Section 3.3, this allows us to easily optimize a variational approximation to the posterior. In addition, it allows for a more complex recurrent dependency on the observed data, via the neural network $f_{p}$, instantiated using a recurrent neural network. The continuous latent variables $z_{t}^{(k)}$ follow an autoregressive model, parametrized by a time-invariant matrix $A$ and a diagonal noise term $\sigma_{p}^{2}$.
[…]

### v2

@@ 1 Introduction
The brain receives information from the surrounding world before reacting correspondingly and systematically. As a complex system, it acts differently across time and locations. Thus, analytical tools that can detect correlations in spatial and temporal activity are required to study the brain’s activity. The dynamic networks that underlie these brain processes at the microscale can be studied using microelectrode array (MEA) recordings of neural signals across a whole population of neural cells (Schröter et al.,, 2017). While MEA recordings do not allow us to target specific neurons, neural signals, highly correlated in time, from neurons near multiple electrodes can reveal the spatial-temporal features of brain network function. Understanding this functional connectivity in vitro is important for understanding microscale brain function and development in vivo (Humphries,, 2017).
@@
Previous studies of microscale functional neuronal networks have focused on the network topology (Downes et al.,, 2012; Schroeter et al.,, 2015; Sit et al.,, 2024). These studies estimate a functional graph by considering pair-wise correlations between electrodes over a period of time. Graph theoretical metrics are then used to compare the topology of the functional networks including the spatial arrangement of nodes (Schröter et al.,, 2017). These approaches have two important limitations. Firstly, they estimate a single pattern of functional connectivity based on minutes of MEA recording. This ignores the fact that brain networks are adaptable on much faster time-scales and precludes study of such short-term changes in connectivity. Secondly, they rely on pair-wise comparison of nodal activity to determine functional connectivity, which may not capture patterns of activity observed in subnetworks detected by more than two electrodes.
@@
Other works have used discrete state-space models to capture evolving functional connectivity. These approaches capture the idea that a dynamical system—in our case, temporal MEA recordings—can be explained in terms of a small number of underlying behavioral patterns. In the simplest state-space model, hidden Markov models (HMMs) model neuronal activity as arising from one of a small number of underlying activity patterns (Paninski et al.,, 2010). Switching linear dynamical systems (SLDS) extend this by characterizing each activity pattern using a linear dynamical system (Petreska et al.,, 2011). Extensions add additional structure between the latent states, allowing for more flexible dynamics (Linderman et al.,, 2017; Nassar et al.,, 2019; Geadah et al.,, 2024; Fox et al.,, 2008). These approaches, however, make the unrealistic assumption that neuronal activity can be partitioned into a small number of a priori independent activity patterns. They lack the capacity to detect multiple network patterns simultaneously, as occurs in complex brain networks.
New methods are needed to identify changes in functional connectivity across seconds to minutes during MEA recordings, without making unrealistic assumptions about the nature of connectivity changes. In this work, we posit that neuronal activity can be decomposed into multiple subnetworks that can be activated jointly or separately. This allows us to capture overlapping patterns of behavior without the need to create a single state for each combination of active subnetworks.
@@
We begin by introducing the relaxed factorial switching model (RFSM), a relaxation of the factorial hidden Markov model (FHMM, Ghahramani and Jordan,, 1995). Like the FHMM, the RFSM allows multiple model components to be turned on and off independently. However, unlike the FHMM, the RFSM is a continuous distribution, allowing us to perform efficient gradient-based variational inference, and allows us to easily relax the Markovian assumption if necessary. We use the RFSM as the backbone of our proposed Factorial Switching Linear Dynamical System (FSLDS). The latent features in the RFSM correspond to subnetworks of neurons, which can be activated independently. When active, each subnetwork’s activity is governed by a linear dynamical system. To facilitate inference, we formulate our latent state model using a continuous relaxation of the factorial hidden Markov model (Maddison et al., 2017b, ) and perform inference using autoencoded variational Bayes (AEVB, Kingma and Welling,, 2014). We first explore properties of FSLDS on simulated data, validating its ability to recover ground-truth additive latent structure. Second, we show how it can uncover interpretable latent structure in MEA recordings. Third, we show how it reveals a developmental increase in the number of latent features (or subnetworks) in human induced pluripotent stem cell (iPSC)-derived neuronal networks in vitro.
@@ 2 Background
Our model is a factorial state-space model that extends existing time-series models by allowing multiple latent factors to be active at once, using a continuous relaxation of a factorial hidden Markov model (FHMM, Ghahramani and Jordan,, 1995) to control which combination of latent factors are active. This leads to a more interpretable model for temporally evolving neuronal functional connectivity, where we assume that observations are generated from an additive combination of multiple subnetworks, that may not all activate synchronously. Optionally, we relax the Markovian assumption to allow more flexible dynamics such as periodicity.
In Section 2.1, we review existing work in state space models for time series data, with a particular focus on models appropriate for modeling MEA recordings. In Section 2.1.1, we focus specifically on factorial state space models, particularly the FHMM which forms the backbone of our approach. We then discuss Bayesian inference in complex discrete state-space models in Section 2.2.
@@ 2.1 Modeling functional connectivity using discrete state space models
Discrete state-space models such as hidden Markov models  (HMMs, Baum and Petrie,, 1966) provide an interpretable model for time-series data, where we assume that each time-point belongs to one of a finite number of states. Observations are assumed iid given a state-specific emission-distribution. HMMs have been used to uncover latent structure in neuronal activity (Camproux et al.,, 1996; Xydas et al.,, 2011; Li et al.,, 2018). Switching dynamical systems replace the emission model with a dynamical system. The simplest example of this is a switching linear dynamical system  (SLDS, Ackerson and Fu,, 1970), where a Markovian discrete state process switches between autoregressive models with state-specific parameters, which can lead to a more flexible model for MEA recordings (Petreska et al.,, 2011). Building on SLDS, more sophisticated models have been developed to handle more complex scenarios, such as recurrent switching linear dynamical system  (rSLDS, Linderman et al.,, 2017) and tree-structured rSLDS Nassar et al., (2019). These models introduce recurrent dependency structures, allowing memory between consecutive time points. Additionally, when linear assumptions prove insufficient, nonlinear dynamical system methods have been proposed Dong et al., (2020); Smith et al., (2021)Ojeda et al., (2021) introduces a variational RNN model with attention to incorporate both nonlinear dynamics and non-Markovian switches.
2.1.1 Factorially structured discrete state space models
$h^{(k)}_{t-1}$$h^{(k)}_{t}$$h^{(k)}_{t+1}$$y_{t-1}$$y_{t}$$y_{t+1}$ $K$
@@ 2.1.1 Factorially structured discrete state space models
The above models assume the underlying behavior switches between a small set of patterns. This may be unrealistic in complex time series such as MEA recordings.
In practice, two timepoints may share some characteristics, but not others. This observation motivates the factorial hidden Markov model (FHMM, Ghahramani and Jordan,, 1995), where multiple hidden factors $\{h_{t}^{(k)}\}_{k=1}^{K}$ concurrently influence the observed dynamics at each time point $y_{t}$.
@@
This factorial structure has previously been used in the context of dynamical systems. Quinn et al., (2008) assume a known switching pattern and use a FHMM to additively modify a single shared dynamical system. Mudrik et al., (2024) use a similar factorial decomposition to combine dynamical systems; however, they use a sparse coding approach rather than explicitly incorporating temporal dynamics in the latent factors. These works focus on understanding and disentangling latent dynamics. Conversely, our approach disentangles the observation space, explaining observations in terms of multiple, additively combining subnetworks.
2.2 Bayesian inference for discrete state space models
The posterior distributions over the parameters of discrete state space models are often learned through Markov chain Monte Carlo (MCMC) using forward filtering backward sampling (West,, 2013; Godsill et al.,, 2004; Olsson and Ryden,, 2011). However, as model complexity increases, MCMC methods can suffer from slow mixing and high computational cost. Recent advances in variational inference provide a more computationally efficient alternative.
Auto-Encoding Variational Bayes  (AEVB, Kingma and Welling,, 2014) offers a flexible “black box” variational inference framework by using neural networks to specify a variational approximation and using gradient descent to optimize a sample-based estimate of the ELBO. However, gradient-based optimization methods cannot directly learn discrete variables, such as those present in switching dynamical systems. The REINFORCE algorithm (Weber et al.,, 2015) uses a Monte Carlo estimate of the gradient; however this estimate suffers from high variance (Paisley et al.,, 2012). Other approaches combine variational inference with sequential Monte Carlo (Maddison et al., 2017a, ; Naesseth et al.,, 2018; Le et al.,, 2018; Lawson et al.,, 2022).
An alternative approach is to sidestep the use of discrete distributions altogether. The Concrete, or Gumbel-Softmax, distribution (Jang et al.,, 2017; Maddison et al., 2017b, ),
$\text{Concrete}(x;\alpha,\phi)=(K-1)!\phi^{K-1}\prod_{k=1}^{K}\left(\frac{\alpha_{k}x_{k}^{-\phi-1}}{\sum_{j=1}^{K}\alpha_{j}x_{j}^{-\phi}}\right)$ (2)
is a continuous distribution over the probability simplex that can be used as an approximation to a Categorical($\alpha_{1},\dots,\alpha_{K}$) distribution. The parameters $\alpha_{1},\dots,\alpha_{K}$ in Equation 2 directly correspond to the probabilities in a Categorical distribution. The temperature $\phi\geq 0$ controls the sparsity of samples. When $\phi=0$, we recover the Categorical($\alpha_{1},\dots,\alpha_{K}$), with samples lying at the vertices of the simplex. As $\phi$ increases, the probability mass shifts towards the center of the simplex. Samples no longer put all their probability mass on a single category. Instead, they assign probability mass to all categories, with samples moving (in expectation) closer to the mean of the Categorical distribution as $\phi$ increases. Since the Concrete($\alpha,\phi$) distribution is continuous, we can use it directly in variational inference as an alternative to the Categorical distribution.
@@ 2.2 Bayesian inference for discrete state space models
3.1 Relaxed factorial switching model
The factorial nature of the FHMM leads to an exponentially growing parameter space. Exploration of this discrete state space is challenging, and traditional MCMC methods are prone to difficult and slow mixing (Ghahramani and Jordan,, 1995). As we discussed in Section 2.2, variational inference provides an appealing alternative to MCMC, allowing us to exploit optimization techniques to learn an approximate posterior. However, as we saw, a naive implementation of variational inference methods does not allow for discrete latent states, since gradient descent is inherently designed for continuous variables.
To overcome this limitation, we introduce a continuous variant of the FHMM by relaxing the discrete latent state indicators into continuous variables using the Concrete distribution (Equation 2). We refer to this model as a relaxed factorial switching model (RFSM). We assume $h_{t}^{(k)}\in[0,1]$ and an additive emission model $f_{\theta}$, so that $h_{t}^{(k)}$ can be seen as turning on or off the $k$th latent feature.
$h_{t}^{(k)}|h_{t-1}^{(k)}\sim$ $\text{Concrete}\left(f_{p}\left(h_{<t}^{(k)},y_{<t}\right),\phi\right)$ (3)
@@ 3.1 Relaxed factorial switching model
In a FHMM, each latent factor is associated with a transition matrix that describes the probability of transitioning between states. In the RFSM, we no longer have discrete state indicators (as $h_{t}^{(k)}$ is continuous), so it no longer makes sense to associate a transition distribution with each latent state. Instead, we specify our transition probabilities using the output of a function $f_{p}$. We have significant flexibility in specifying $f_{p}$. We can construct a Markovian switching model by specifying $f_{p}$ using a multi-layer perceptron (MLP) that takes as inputs $h_{t-1}$ and $y_{t-1}$. Or, we can relax the Markovian assumption and use a non-Markovian function such as a recurrent neural network (RNN). In our experiments, we explore both approaches. In datasets where we observe stereotypy, we use RNNs; in datasets with no obvious stereotypy, we use the simpler MLP.
@@ 3.2 Factorial switching linear dynamical systems
As we have seen, HMMs, SLDSs, and their variants assume each time point corresponds to a single latent state. In practice, this might be unrealistic. For example, in developing brains, we expect new subnetworks to appear in addition to previous structure, rather than fully replacing the previous connectivity pattern. Further, we do not expect a subnetwork to have constant activity—instead, we want to disentangle changes in activity amplitude from changes in neuron connectivity. To achieve this, we first specify a set of $K$ features $\theta_{k}\in\mathbb{R}_{+}^{M}$, where $\theta_{k,m}$ represents the relative contribution [^1] Since we scale $\theta_{k}$ by a continuous rate, $\theta_{k,m}$ can only be interpreted as relative contribution of the $m$th node relative to other nodes. of the $m$th node to the $k$th subnetwork. We then use an RFSM to determine whether that feature is active at a given time, and model the value of the state when it is active using a linear dynamical system. The generative model is
@@
$h_{t}^{(k)}\mid h_{t-1}^{(k)}$ $\sim\text{Concrete}\left(f_{p}\left(h_{t-1}^{(k)},\,y_{t-1}\right),\phi\right),$ $k=1,\dots,K$
$z_{t}^{(k)}\mid z_{t-1}^{(k)}$ $\sim\text{Normal}\left(A^{(k)}z_{t-1}^{(k)},\,\sigma_{p}^{2}\mathbb{I}\right),$ $k=0,\dots,K.$
The level of activity of the $k$th subnetwork $\theta_{k}$ at time $t$ is governed by two components: an (approximately) discrete state $h_{t}^{(k)}\in[0,1]$, and a continuous state $z_{t}^{(k)}\in\mathbb{R}$. The approximately discrete state $h_{t}^{(k)}$ controls whether a subnetwork is currently contributing to the rate, and aids in interpretability of the latent structure. Meanwhile the continuous states $z_{t}^{(k)}$ control variation in the amplitude of the activity. We include an “always-on” discrete background state $h_{t}^{(0)}=1$, and associated continuous state $z_{t}^{(0)}$, to capture persistent background noise. In the context of MEA recordings, we can think of $h_{t}^{(k)}$ as indicating whether a subnetwork is currently active, and $\exp\left\{z_{t}^{(k)}\right\}$ as controlling variation in its activity while it is active. We summarize FSLDS in Figure 2.
The choice of an RFSM to model $h_{t}^{(k)}$ is primarily due to computational practicality; as we will see in Section 3.3, this allows us to easily optimize a variational approximation to the posterior. In addition, if we have non-Markovian switching patterns, it allows for a more complex recurrent dependency on the observed data (by instantiating $f_{p}$ using an RNN). The continuous latent variables $z_{t}^{(k)}$ follow an autoregressive model, parametrized by a time-invariant matrix $A^{(k)}$ and a diagonal noise term $\sigma_{p}^{2}$.
$h^{(k)}_{t-1}$$h^{(k)}_{t}$$h^{(k)}_{t+1}$$z^{(k)}_{t-1}$$z^{(k)}_{t}$$z^{(k)}_{t+1}$$y_{t-1}$$y_{t}$$y_{t+1}$$\theta^{(k)}$ $K$ $K$ $K$